# 22-descent for Bloch–Kato Selmer groups and rational points on hyperelliptic curves I

### v3

@@ 5. $2$-descent for Selmer schemes: local aspects at 22
More precisely, let $b$ and $z$ be $\mathbb{Z}_{2}$ points of $X$ not reducing to the point at infinity, and let $\Pi_{2}$ denote the maximal $2$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{\mathbb{Q}}_{2}}-\{\infty\},b)$. Then $j_{\Pi_{2},b}(z)$ is an element of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$, i.e. an element of $H^{1}(\mathbb{Q}_{2},\Pi_{2})$ whose image in $H^{1}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ lies in the image of the subset $H^{1}_{f}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ in the sense of Kim [Kim09]. Furthermore, if zbz-b lies in $4\cdot J(\mathbb{Q}_{2})$, then $j_{\Gamma,b}(z)$ lifts to an element of $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, which we shall also denote $j_{\Gamma,b}(z)$ (although if $J[4](\mathbb{Q}_{2})\neq 0$ then this lift is not unique).
@@ 5.1. Integral mixed extensions
$T_{0}^{*}\otimes T_{1}\oplus T_{1}^{*}\otimes T_{2}$
@@
$(g,h)\mapsto\left(\begin{array}[]{ccc}1&0&0\\ 0&0&0\\ \beta(g)\cdot(g\cdot\alpha(h))&0&0\end{array}\right),$
@@
If $M$ is an integral mixed extension with graded pieces $T_{0},T_{1},T_{2}$, such that $M/T_{2}]\in p\cdot H^{1}_{f}(\mathbb{Q}_{p},T_{0}^{*}\otimes T_{1})$, then there is a Galois stable sublattice $M^{\prime}$ of $M$ with $M/M^{\prime}\simeq T_{1}\otimes\mathbb{F}_{p}$, and such that
@@
Let $\theta:H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)\times H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)\to H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,T_{J}^{\otimes 2}))$ be the map induced by the left and right multiplication maps
@@
$\widehat{J}^{\otimes}2\to\widehat{J}^{\otimes 2}$
@@ 6.1. Galois cohomology of number fields
Let $S_{0}$ be a finite set of primes, and suppose $\mathcal{X}$ is a proper regular model of $X$ over $\mathbb{Z}_{S_{0}}$. Then the étale regulator
@@
    [FPR94, Remark 1.2.4]:If $K=\mathbb{Q}$, and $H^{0}(G_{\mathbb{Q},T},W)$, $H^{0}(G_{\mathbb{Q},T},W^{*}(1))$ and $D_{\mathrm{cris}}(W)^{\phi=1}$ are all zero, then
@@
    [FPR94, 1.2.2]For any number field $K$ and $\mathbb{Q}_{p}$-Galois representation $W$,
@@ 6.2. The Chabauty–Coleman–Kim sets $X(\mathbb{Q}_{p})_{n}$
Let $X$ be a smooth projective geometrically irreducible curve over $\mathbb{Q}$, of genus $g>1$. Let $b\in X(\mathbb{Q})$. Fix a prime $p$. The Chabauty–Coleman–Kim method, as developed in [Kim09], [KT08] and [Kim05], determines subsets $X(\mathbb{Q}_{p})_{n}\subset X(\mathbb{Q}_{p})$, containing the set of rational points $X(\mathbb{Q})$, which are defined in terms of the $\mathbb{Q}_{p}$-unipotent fundamental group of $X$, which we now define.
@@ 6.3. Dimension bounds via descent
For any finite set $S$ of primes containing all archimedean places, we have a short exact sequence
@@
We have
@@
7. Computation of the generalised height
We now explain how to compute $X(\mathbb{Q}_{2})_{2}$ using the ‘quadratic’ structure of $H^{1}(K,U_{2})$, following [BD21]. Here is a heuristic explanation. For simplicity we assume that $X$ is a hyperelliptic curve with a chosen rational Weierstrass point $\infty$. Mimicking the construction in the ‘usual’ version of quadratic Chabauty, we seek functions
$h_{v}:X(\mathbb{Q}_{v})\to\mathbb{Q}_{p}$
such that we can prove there is a linear map
$h:J(\mathbb{Q})^{\otimes 2}\to\mathbb{Q}_{p},$
with the property that when $x$ lies in $X(\mathbb{Q})$, we have
$\sum h_{v}(x)=h(\mathrm{AJ}(x-\{\infty\}),\mathrm{AJ}(x-\{\infty\})).$
Strictly speaking, in general we construct slightly less than this, as in general this linear map is only partially defined. For this reason, it is better to say that the functions $h_{v}$ should satisfy the property that if $x_{i}\in X(\mathbb{Q})$, for $i=1,\ldots,n$, and $\sum n_{i}[x_{i}]^{2}=0$ in $J(\mathbb{Q})^{\otimes 2}$, then
(18) $\sum_{v}\sum_{i}n_{i}h_{v}(x_{i})=0.$
In fact, there will be many such height functions $h_{v}$ that we want to use, and it is better to linearly combine them all to construct vector space-valued heights: specifically we will construct local heights with values in $H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$, so that the identity (18) is taking place in
$\oplus_{v\in S}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1)).$
The reader familiar with the theory of $p$-adic height pairings over number fields may think of $\oplus_{v\in S}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ as a generalisation of the space of functionals on idele class characters.
We first recall the definition and basic properties. We refer the reader to [BD21, §3] for more details.
Lemma 28 ([BD21], §3).
For $v\neq p$, the map
$H^{1}(\mathbb{Q}_{v},\wedge^{2}V/\mathbb{Q}_{p}(1))\to H^{1}(\mathbb{Q}_{v},U_{2})$
induced by the inclusion of groups $\wedge^{2}V/\mathbb{Q}_{p}(1)\to U_{2}$ is an isomorphism.
We define
$\widetilde{h}_{v}:H^{1}(\mathbb{Q}_{v},U_{2})\to H^{1}(\mathbb{Q}_{v},\wedge^{2}V/\mathbb{Q}_{p}(1)),$
to be the inverse of this isomorphism. For $v=p$, the definition of the local height
$\widetilde{h}_{p}:H^{1}_{f}(\mathbb{Q}_{p},U_{2})\to H^{1}_{f}(\mathbb{Q}_{p},\wedge^{2}V/\mathbb{Q}_{p}(1))$
is recalled in the next section.
Proposition 5 ([BD21], §3).
Let $X$ be a hyperelliptic curve with hyperelliptic involution $w$. Then the local heights $\widetilde{h}_{p}$ have the property that if
$\sum n_{i}(P_{i}-b_{i})\otimes(P_{i}-w(b_{i}))=0$
in $J(\mathbb{Q})^{\otimes 2}\otimes\mathbb{Q}$, then there is an element $\xi\in H^{1}_{f,S_{0}}(\mathbb{Q},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ such that
$\sum n_{i}\widetilde{h}_{v}(j_{v,b_{i}}(P_{i}))=\mathrm{loc}_{v}(\xi)$
for all $v\in S$.
It follows that the map
$\mathbb{Q}_{p}[X(\mathbb{Q})]\to\oplus_{v}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$
sending $z$ to $\sum_{v}\widetilde{h}_{v}(j_{v,\infty}(z)$ factors through the map
$\mathbb{Q}_{p}[X(\mathbb{Q})]\to J(\mathbb{Q})^{\otimes 2}\otimes\mathbb{Q}_{p}.$
To get a symmetric map (i.e. a map factoring through $\mathrm{Sym}^{2}J(\mathbb{Q})$) we can remove the alternating part as follows. First recall (from [BD21]) that the generalised height is really a map on the set $M(\mathbb{Q}_{p},V_{p}J,\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ of equivalence classes of mixed extensions of $\mathbb{Q}_{p}$-Galois representations with graded pieces $\mathbb{Q}_{p},V_{p}J$ and $\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1)$. The map on $X(K)$ (for $K=\mathbb{Q}$ or $\mathbb{Q}_{v}$) is then defined by sending $x$ to the generalised height evaluated at the mixed extension $E(b,x)_{\mathbb{Q}_{2}}$, where $E(b,x)_{\mathbb{Q}_{2}}$ is defined as follows. Let $E_{n}(b,x)$ be the Galois representation defined in section 5. Let $E_{n}(b,x)_{\mathbb{Q}_{2}}:=E_{n}(b,x)\otimes_{\mathbb{Z}_{2}}\mathbb{Q}_{2}$. Then $E_{2}(b,x)_{\mathbb{Q}_{2}}$ is an extension of $E_{1}(b,x)_{\mathbb{Q}_{2}}$ by
$\mathrm{Ker}(H^{1}_{\mathrm{\acute{e}t}}(X_{\overline{K}},\mathbb{Q}_{p})^{\otimes 2}\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}H^{2}_{\mathrm{\acute{e}t}}(X_{\overline{K}},\mathbb{Q}_{p}))^{*}.$
Define $E(b,x)$ to be $E_{2}(b,x)_{\mathbb{Q}_{2}}/\mathrm{Sym}^{2}V_{p}(J)$.
Using the notion of equivariant generalised heights (see [BD21, §4]) one can also evaluate such a generalised height on a mixed extension with graded pieces $\mathbb{Q}_{p},V_{p}J^{\oplus n}$ and $\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1)$. Now for $M_{1},M_{2}\in M(\mathbb{Q}_{p},V)$, define $M_{1}\wedge M_{2}\in M(\mathbb{Q}_{p},V^{\oplus 2},\wedge^{2}V/\mathbb{Q}_{p}(1))$ to be the quotient of $M_{1}\otimes M_{2}$ by $\mathrm{Sym}^{2}V_{p}J\oplus\mathbb{Q}_{p}(1)$. We define
$h_{v}(z_{1},z_{2}):=\widetilde{h}_{v}(E(z_{1},z_{2}))-\frac{1}{2}\widetilde{h}_{v}(E_{1}(z_{1},z_{2})_{\mathbb{Q}_{2}}\wedge E_{1}(z_{2},w(z_{2}))_{\mathbb{Q}_{2}}).$
We deduce the following.
Lemma 29.
The map
$h:\mathbb{Q}_{p}[X(\mathbb{Q})\times X(\mathbb{Q})]\to\oplus_{v}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$
factors through the map
$\mathbb{Q}_{p}[X(\mathbb{Q})\times X(\mathbb{Q})]\to\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}_{p}$
sending $(z_{1},z_{2})$ to $(z_{1}-z_{2})\cdot(z_{1}-w(z_{2}))$.
Note that when $v\neq p$, $h_{v}=\widetilde{h}_{v}$ since $H^{1}(\mathbb{Q}_{v},V)=0$. The question of how to compute the local heights $\widetilde{h}_{v}$ for $v\neq p$ is a special case of the question of how to describe the action of monodromy on the unipotent fundamental group and associated path torsors. This was studied by Kim and Tamagawa [KT08] and also in recent work with Betts [BD19]. For studying the BMSST curve, we will only need the following result.
Lemma 30 ([BD19], Proposition 3.8.1).
Suppose that $x_{1},x_{2}\in X(\mathbb{Q}_{v})$ lie on a common irreducible component are regular semistable model of $X$ over a finite extension of $\mathbb{Q}_{v}$. Then the class of $x_{2}$ in $H^{1}(G_{\mathbb{Q}_{v}},U_{n}(x_{1}))$ is trivial for all $n$.
In particular, together with Lemma 28 this implies that the local height of $E_{2}(x_{1},x_{2})$ at $v$ will be trivial.
7.1. Computing $X(\mathbb{Q}_{2})_{2}$
We compute the set $X(\mathbb{Q}_{2})_{2}$ as follows. Let
$h_{2}:X(\mathbb{Q}_{2})\to\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*}/\langle H^{2}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{*},\wedge^{2}(F^{1}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{2})^{\perp}\rangle$
denote the generalised height with respect to a basepoint $b$. For $v\neq 2$, let $\Omega_{v}\subset H^{1}(\mathbb{Q}_{v},\wedge^{2}T_{2}V/\mathbb{Q}_{p}(1))$ denote the image of $X(\mathbb{Q}_{v})$ under $j_{v}$. Suppose for simplicity that all the $\alpha_{v}$ are trivial (this is the setting we will be in in the case of the BMSST curve). Let $X$ be a hyperelliptic curve of genus $g$ and Mordell–Weil rank $g\leq r<g^{2}$. Suppose that we have verified the rank part of the Bloch–Kato conjecture for $H^{1}_{f}(\mathbb{Q},\wedge^{2}V_{p}J)$, so that $\dim H^{1}_{f}(\mathbb{Q},\wedge^{2}V_{p}J)=(g-1)(g+2)/2$ and $X(\mathbb{Q}_{p})_{2}$ is finite. For a divisor $D=\sum n_{i}z_{i}$, we will slightly bizarrely write $h_{v}(D)$ to mean $\sum_{i}n_{i}h_{v}(z_{i})$. Define $\Theta$ to be the map
$\mathrm{Div}(X)(\mathbb{Q})\stackrel{{\scriptstyle(\mathrm{AJ}_{\mathbb{Q}_{p}},\mathrm{Sym}^{2}\mathrm{AJ}_{\mathbb{Q}_{p}},h_{p}(E))}}{{\longrightarrow}}J(\mathbb{Q})\otimes\mathbb{Q}_{p}\oplus\mathrm{Sym}^{2}J(\mathbb{Q})\otimes\mathbb{Q}_{p}\oplus\mathrm{Fil}^{0}H^{2}_{\mathrm{dR}}(J_{\mathbb{Q}_{p}}/\mathbb{Q}_{p})^{*}$
Then the generalised height formalism, together with the verification of the rank part of the Bloch–Kato conjectures, tells us that $\Theta$ maps into an N:=r(r+3)/2+(g1)(g+2)/2N:=r(r+3)/2+(g-1)(g+2)/2-dimensional subspace of the target, which is an $N+g^{2}-g$ dimensional vector space.
Suppose we have divisors $D_{1},\ldots,D_{N}$ such that $\langle\Theta(D_{1}),\ldots,\Theta(D_{N})\rangle$ has dimension $N$. Then for any point $z\in X(\mathbb{Q})$, we have the nontrivial relation
$\operatorname{rk}(\Theta(D_{1}),\ldots,\Theta(D_{N}),\Theta(z))=N.$
This is not yet an identity which can be tested on $p$-adic points, because it involves the image of zbz-b in $J(\mathbb{Q})\otimes\mathbb{Q}_{p}$, which does not inject into $H^{0}(X_{\mathbb{Q}_{p}},\mathbb{Q}_{p})^{*}$. We can remedy this by elimination of variables.
Suppose that $J(\mathbb{Q})$ generated $H^{0}(X_{\mathbb{Q}_{p}},\Omega)^{*}$ as a $\mathbb{Q}_{p}$-vector space, and define functionals $\psi_{1},\ldots,\psi_{r}:J(\mathbb{Q})\to\mathbb{Q}_{p}$ linearly independent from $\log_{J}$. Hence we may choose our basis of $J(\mathbb{Q})\otimes\mathbb{Q}_{p}$ so that $\mathrm{AJ}_{Q_{p}}$ is given by $(\log_{J},\psi_{1},\ldots,\psi_{r})$.
Define $\Theta_{p}:X(\mathbb{Q}_{p})\to\mathbb{Q}_{p}[t_{1},\ldots t_{r-g}]^{N}$ to be the map
$(\log_{J},t_{1},\ldots,t_{r},\mathrm{Sym}^{2}(\log_{J},t_{1},\ldots,t_{r}),h_{p}(E(.))).$
Then we deduce the following.
Lemma 31.
The set $X(\mathbb{Q}_{p})_{2}\subset X(\mathbb{Q}_{p})$ is contained in the set of $z\in X(\mathbb{Q}_{p})$ such that
$\operatorname{rk}(\Theta_{p}(D_{1})(\boldsymbol{\psi}(D_{1})),\ldots,\Theta_{p}(D_{N})(\boldsymbol{\psi}(D_{N})),\Theta_{p}(z))=N$
has a solution in $\mathbb{Q}_{p}^{r-g}$.
8. $2$-adic Coleman integration
The computation of the local height at $p$ involves calculating Frobenius structures and Hodge filtrations on certain filtered $F$-isocrystals $\mathcal{E}$ on $X$ (see [BDM+19, §4] for a similar discussion in the context of ‘non-generalised’ quadratic Chabauty). Roughly speaking, following Nekovar, in [BD21] a local $p$-adic height at $p$ is defined to be a function on certain isomorphism classes of filtered $\phi$-modules (with extra structure). These filtered $\phi$-modules are the fibres of $\mathcal{E}$ at $\mathbb{Z}_{p}$-points of $X$.
As explained above, the $2$-descent methods we use can only prove finiteness of $X(\mathbb{Q}_{2})_{2}$, not of $X(\mathbb{Q}_{p})_{2}$ for $p>2$. Hence, to apply them in examples, we need to compute $2$-adic local heights at $2$, which in particular involves computing $2$-adic Coleman integrals. The computation of Coleman integration may be described as having two aspects. The first, easier aspect is that of ‘tiny integrals’. This is essentially the same for all primes $p$ of good reduction (and even primes of bad reduction, once the definitions are suitably interpreted).
The second, more complicated aspect, is that of Frobenius structures on unipotent connections. As we recall below, a Frobenius structure on an isocrystal is an isomorphism between the isocrystal and its pullback by Frobenius. Hence the computation of a Frobenius structure itself has two elements: first the computation of a lift of Frobenius, and second the computation of the isomorphism of isocrystals. In practice, for unipotent isocrystals, an algorithm for computing this isomorphism essentially amounts to an algorithm for presenting any $n$-unipotent isocrystal as a quotient of a certain ‘universal’ $n$-unipotent isocrystal. This problem essentially amounts to a reduction algorithm for overconvergent differentials.
For hyperelliptic curves over $p$-adic fields when $p>2$, the problem of algorithmically lifting Frobenius and reducing differentials was solved by Kedlaya [Ked01]. The problem of generalising Kedlaya’s algorithm to the case $p=2$ was considered by Denef and Vercauteren [DV02] [DV06]. What we do below is a particularly simple case of the problem of extending their work to give an algorithm for computing Frobenius structures on unipotent isocrystals on Artin–Schreier or hyperelliptic curves in characteristic 2.
Namely, in this section, we consider a slightly less general situation than theirs, which includes the BMSST curve. We consider a smooth hyperelliptic curve over $\mathbb{Z}_{2}$ of the form
$y^{2}-y=f(x),$
where $f(x)=\sum_{i=0}^{2g+1}a_{i}x^{i}$ has degree $2g+1$. Let $Q(x,y):=y^{2}-y-f(x)$. Define $(y_{n})$ by
$y_{1}$ $=y^{2}$
$y_{n+1}$ $=y_{n}-Q(x^{2},y_{n}).$
Lemma 32.

### v4

@@ 5. $2$-descent for Selmer schemes: local aspects at 22
More precisely, let $b$ and $z$ be $\mathbb{Z}_{2}$ points of $X$ not reducing to the point at infinity, and let $\Pi_{2}$ denote the maximal $2$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{\mathbb{Q}}_{2}}-\{\infty\},b)$. Then $j_{\Pi_{2},b}(z)$ is an element of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$, i.e. an element of $H^{1}(\mathbb{Q}_{2},\Pi_{2})$ whose image in $H^{1}(\mathbb{Q}_{2},U_{2})$ lies in the subset $H^{1}_{f}(\mathbb{Q}_{2},U_{2})$ in the sense of Kim [Kim09]. Here $U_{2}$ denote the maximal $2$-nilpotent quotient of the $\mathbb{Q}_{p}$-unipotent fundamental group of $X_{\overline{\mathbb{Q}}_{2}}$-$\infty$. Furthermore, if zbz-b lies in $4\cdot J(\mathbb{Q}_{2})$, then $j_{\Gamma,b}(z)$ lifts to an element of $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, which we shall also denote $j_{\Gamma,b}(z)$ (although if $J[4](\mathbb{Q}_{2})\neq 0$ then this lift is not unique).
@@ 5.1. Integral mixed extensions
$U(T_{0},T_{1},T_{2})\to T_{0}^{*}\otimes T_{1}\oplus T_{1}^{*}\otimes T_{2}$
@@
$(g,h)\mapsto\beta(g)\cdot(g\cdot\alpha(h)),$
@@
If $M$ is an integral crystalline mixed extension with graded pieces $T_{0},T_{1},T_{2}$, such that $M/T_{2}]\in p\cdot H^{1}_{f}(\mathbb{Q}_{p},T_{0}^{*}\otimes T_{1})$, then there is a Galois stable sublattice $M^{\prime}$ of $M$ with $M/M^{\prime}\simeq T_{1}\otimes\mathbb{F}_{p}$, and such that
@@
Let $\theta:H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)\times H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)\to H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,T_{2}J^{\otimes 2}))$ be the map induced by the left and right multiplication maps
@@
$\widehat{J}^{\otimes 2}\times\widehat{J}^{\otimes 2}\to\widehat{J}^{\otimes 2}$
@@ 6.1. Galois cohomology of number fields
Let $S_{0}$ be a finite set of primes, and suppose $\mathcal{X}$ is a proper regular model over $\mathbb{Z}_{S_{0}}$ of a smooth projective variety $X/\mathbb{Q}$. Then the étale regulator
@@
Let $W$ be a finite dimensional $\mathbb{Q}_{p}$ representation of $\operatorname{Gal}(K)$ unramified outside a finite set of primes.
@@
    [FPR94, Remark 1.2.4]
    If $K=\mathbb{Q}$, and $H^{0}(G_{\mathbb{Q},T},W)$, $H^{0}(G_{\mathbb{Q},T},W^{*}(1))$ and $D_{\mathrm{cris}}(W)^{\phi=1}$ are all zero, then
@@
    [FPR94, 1.2.2]
    For any number field $K$ and $\mathbb{Q}_{p}$-Galois representation $W$,
@@ 6.2. The Chabauty–Coleman–Kim sets $X(\mathbb{Q}_{p})_{n}$
Let $X$ be a smooth projective geometrically irreducible curve over $\mathbb{Q}$, of genus $g>1$. Let $b\in X(\mathbb{Q})$. Fix a prime $p$ of good reduction for $X$. The Chabauty–Coleman–Kim method, as developed in [Kim09], [KT08] and [Kim05], determines subsets $X(\mathbb{Q}_{p})_{n}\subset X(\mathbb{Q}_{p})$, containing the set of rational points $X(\mathbb{Q})$, which are defined in terms of the $\mathbb{Q}_{p}$-unipotent fundamental group of $X$, which we now define.
@@ 6.3. Dimension bounds via descent
Let $L$ be a number field. For any finite set $S$ of primes containing all archimedean places, we have a short exact sequence
@@
Let $J$ be the Jacobian of a hyperelliptic curve over a number field $K$ with an odd degree model of the form $y^{2}=f(x)$, and Jacobian $J$. We have
@@
6.4. Generalised heights
We briefly describe how to explicitly compute $X(\mathbb{Q}_{2})_{2}$ using the above. More details may be found in [Dog26]. Suppose for simplicitly that $X$ is a smooth hyperelliptic curve of genus $g>1$ with the following properties:
  • $X$ has good reduction at 2.
  • $X$ has a rational Weierstrass point $\infty$.
  • At all primes $v$, all points in $X(\mathbb{Q}_{v})$ map to a common irreducible component of the special fibre of the minimal regular model of at $v$.
Then, as in [BD21], one constructs a function
$h_{2}:X(\mathbb{Q}_{2})\to W:=H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}V/\mathbb{Q}_{2}(1))/\mathrm{loc}_{2}H^{1}_{f}(\mathbb{Q},\wedge^{2}V/\mathbb{Q}_{2}(1))$
with the following properties: